OpenStudy (anonymous):

does the limit of [ (cos x -1) / sin x ] as x ->0 exist?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\cos(x)-1}{\sin(x)} \cdot \frac{\frac{1}{x}}{\frac{1}{x}}\] the next step is the answer

OpenStudy (anonymous):

so does the limit not exist because the right and left side limits are not equal to each other or because they are increasing without bound?

myininaya (myininaya):

limit of top is 0 limit of bottom is 1 0/1=0

OpenStudy (anonymous):

thanks!

myininaya (myininaya):

you should know that \[\lim_{x \rightarrow 0}\frac{\cos(x)-1}{x}=0 \& \lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1\]